using ordered pairs find the missing coordinate to make the slope true
(x,3) and (19,-17)
Slope = -2

Answer by TimothyLamb(4379)

Get the equation in point-slope form: y - y1 = m(x - x1)
---
Use the given slope -2 and the given point (19,-17), like so:
---
y - (-17) = (-2)(x - 19)
Simplify, like so:
y + 17 = -2x + 38
y = -2x + 21
---
Now plug-in the y-coordinate (3) of the point with missing x-coordinate (x,3), and solve the equation for x, like so:
---
3 = -2x + 21
2x = 21 - 3
2x = 18
Answer:
x = 9
---
To check the answer, calculate the slope from the two points, like so:
---
m = (-17 - 3) / (19 - 9)
m = -20 / 10
m = -2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php